Can you please help me solve this problem? Solve for x: Log10(x-2)+log10(x+7)=1
This problem is from college algebra textbook chapter 3


This is what i have so far
log10(x-2)(x-7)=1
(x-2)(x-7)=10
x2-9x+14=10
Thank you

Answer by venugopalramana(3286)

Can you please help me solve this problem? Solve for x: Log10(x-2)+log10(x+7)=1
This problem is from college algebra textbook chapter 3

This is what i have so far
log10(x-2)(x-7)=1.......NO......(X+7)
(x-2)(x+7)=10.......CORRECTED..
x2+5x-14=10 ........CORRECTED
X^2+5X-14-10=0
X^2+5X-24=0
X^2+8X-3X-24=0
X(X+8)-3(X+8)=0
(X-3)(X+8)=0
X=3.SINCE X=-8 LEADS TO LOG OF NEGATIVE NUMBERS..